OpenStudy (anonymous):

how to do this? lim x->0 (x-9)/x^2/(x+4)

OpenStudy (anonymous):

is your question like this \[ \lim_{x \rightarrow 0}\frac{ x-2 }{\frac{ x ^{2} }{ x+4 }}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cmon there must be someone who knows how to do this :S

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

is \[\frac{ x-9 }{ x ^{2} }\] a numerator and x+4 denominator?

OpenStudy (anonymous):

its (x-9)/(x^2 * (x+4))

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

this goes to heaven ... try pluggin in WA.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ x-9 }{ x ^{2}(x+4)}\]

OpenStudy (experimentx):

yes ... the limit tends to infinity. accordinng to Wolf.

OpenStudy (anonymous):

ok it might tends to infinity because i've tried all the method i know to solve but i remain with x at the denominator

OpenStudy (experimentx):

yes ... limit is inifnity .... or undefined or does not exits.

OpenStudy (anonymous):

u r right